OpenStudy (anonymous):

solve for you: 2(u+1)=-7u+11

OpenStudy (vispla22):

you can distribute the first equation into 2u+2

OpenStudy (vispla22):

which will then turn into 2u+2=-7u+11

OpenStudy (vispla22):

adding 7u to both sides you can get 9u+2=11

OpenStudy (vispla22):

minus 2 to both sides and get 9u=9

OpenStudy (vispla22):

so then u=1

OpenStudy (anonymous):

so u=1

OpenStudy (vispla22):

yep

OpenStudy (anonymous):

thankyou

OpenStudy (camerondoherty):

Step 1 : Simplify \[2•(u+1) - -7u+11 \] Pull out like terms : \[ 9u - 9 = 9 • (u - 1) \] \[ 9 • (u - 1) = 0 \] Step 2 : Solve:\[ 9•(u-1) = 0 \] Equations which are never true : Solve : 9 = 0 This equation has no solution. A a non-zero constant never equals zero. Solve : \[ u-1 = 0 \] Add 1 to both sides of the equation : u = 1 Answer: u = 1